# Knowledge Discovery using Unsupervised Cognition

### v1

@@ 3 Pattern Mining
Unsupervised Cognition [9] is an unsupervised learning method that builds representations in a hierarchical structure. To briefly resume its functioning, it builds representations by combining multiple inputs, thus each representation depicting a subset of the input domain. Then, for each representation, it builds children representations that depict subsets of the set depicted by the parent representation, until there is no more possible subsets. Moreover, each input is assigned to a representation based on a similarity measure, thus each representation depicts a cluster of the input domain. This way, the Unsupervised Cognition algorithm builds a tree of representations where the seed node contains the most generic representations, that is, those that depict the biggest clusters of the input domain; and the leaf nodes contain representations that depict only one input, that is, literal representations. Here it is important to remark that representations at the same level of tree depth do not share any input, that is, any input will belong to only one representation at that level.
@@
Our proposal is to extract patterns from the representations built by the Unsupervised Cognition algorithm. To do so, first we need to define what we consider to be a pattern. In our case, we will consider patterns those representations that are the most generic ones. Under no constraints, the most generic representations build by the Unsupervised Cognition algorithm are those present in its seed node, as they are the ones that aggregated the most quantity of inputs. At the same time, these representations have split the input domain into multiple clusters (one per representation) based on similarity distance, thus telling us that there are different forms an input can take. And due to the fact that representations at the same tree depth do not share inputs, then these representations are the different patterns present in the data.
However, if we have some constraints, then the subset of representations that we will consider patterns change. For example, if we have trained an Unsupervised Cognition model with the purpose of solving a classification task, then the representations will include information of the class they belong (even though the learning algorithm is fully unsupervised). In this case, it is common that some representations at multiple levels of the tree belong to more than one class, because they combined inputs belonging to different classes. Thanks to the set - subset relationship of the parent and children representations, we know that those representations that belong to multiple classes are in the top of the tree, while the representations that belong to only one class are at the bottom of the tree. In this situation, we may want to add the constraint that the patterns differentiate between classes. Given the previous organisation, in this case we will consider patterns those representations that are in the limit, that is, those that belong to only one class, but that at the same time are the most generic ones possible. These representations will have parent representations that belong to multiple classes (unless they are in the seed node) and all their children representations (if they have any) will belong to the same class than them. This setup provides us with at least one pattern for each class, and we may have multiple patterns for the same class.
@@ 4 Dimensionality Reduction
With this rationale, we devised the following method: we start by picking the patterns generated by the Unsupervised Cognition algorithm, and we compute, for each feature, the correlation between the values of such feature in the patterns and the value of the target feature in the patterns. If the target feature is a numerical feature, the patterns will be the representations from the seed node of the Unsupervised Cognition model. However, if the target feature is a categorical feature, then the patterns should be those representations that belong to only one class of the classes present in the categorical feature. In this last case, the correlations will be between each class of the target feature and the values in the other feature, thus generating a correlation for each class of the target feature, for each feature of the input domain.

### v2

@@ 3 Pattern Mining
Unsupervised Cognition [9] is an unsupervised learning method that builds representations in a hierarchical structure. To briefly resume its functioning, it builds representations by combining multiple inputs, thus each representation depicting a subset of the input domain. Then, for each representation, it builds children representations that depict subsets of the set depicted by the parent representation, until there is no more possible subsets. Moreover, each input is assigned to a representation based on a similarity measure, thus each representation depicts a cluster of the input domain. This way, the Unsupervised Cognition algorithm builds a tree of representations where the seed Cell contains the most generic representations, that is, those that depict the biggest clusters of the input domain; and the leaf Cells contain representations that depict only one input, that is, literal representations. Here it is important to remark that representations at the same level of tree depth do not share any input, that is, any input will belong to only one representation at that level.
@@
Our proposal is to extract patterns from the representations built by the Unsupervised Cognition algorithm. To do so, first we need to define what we consider to be a pattern. In our case, we will consider patterns those representations that are the most generic ones. Under no constraints, the most generic representations build by the Unsupervised Cognition algorithm are those present in its seed Cell, as they are the ones that aggregated the most quantity of inputs. At the same time, these representations have split the input domain into multiple clusters (one per representation) based on similarity distance, thus telling us that there are different forms an input can take. And due to the fact that representations at the same tree depth do not share inputs, then these representations are the different patterns present in the data.
However, if we have some constraints, then the subset of representations that we will consider patterns change. For example, if we have trained an Unsupervised Cognition model with the purpose of solving a classification task, then the representations will include information of the class they belong (even though the learning algorithm is fully unsupervised). In this case, it is common that some representations at multiple levels of the tree belong to more than one class, because they combined inputs belonging to different classes. Thanks to the set - subset relationship of the parent and children representations, we know that those representations that belong to multiple classes are in the top of the tree, while the representations that belong to only one class are at the bottom of the tree. In this situation, we may want to add the constraint that the patterns differentiate between classes. Given the previous organisation, in this case we will consider patterns those representations that are in the limit, that is, those that belong to only one class, but that at the same time are the most generic ones possible. These representations will have parent representations that belong to multiple classes (unless they are in the seed Cell) and all their children representations (if they have any) will belong to the same class than them. This setup provides us with at least one pattern for each class, and we may have multiple patterns for the same class.
@@ 4 Dimensionality Reduction
With this rationale, we devised the following method: we start by picking the patterns generated by the Unsupervised Cognition algorithm, and we compute, for each feature, the correlation between the values of such feature in the patterns and the value of the target feature in the patterns. If the target feature is a numerical feature, the patterns will be the representations from the seed Cell of the Unsupervised Cognition model. However, if the target feature is a categorical feature, then the patterns should be those representations that belong to only one class of the classes present in the categorical feature. In this last case, the correlations will be between each class of the target feature and the values in the other feature, thus generating a correlation for each class of the target feature, for each feature of the input domain.